i have sales data for the period jan to june 2016

how do i calculate the implied seasonal index for each month

Actual Seasonally adjusted index

jan 150 176 ?

feb 140 175 ?

mar 130 173 ?

apr 120 185 ?

may 110 183 ?

june 190 158 ?

the trend equation based on the last 6 months seasonally adjusted data is ?

seasonally adjusted sale Y = 150 + 10k so than do a forecast of actual month sale for next june 2017

time value (x) for next june is 10 and use the seasonal index from last june 2016
Answer #1

For ease of understanding I have named the given months from Jan15-Jun16. Also, we are considering that there are only 6 months in a year. We will be calculating the seasonality of each month (ie Jan – Jun) and use it to predict Sales for next 6 months (Jan17-Jun17)

Steps:

  1. First we calculate 4 months moving Average.
  2. Then we calculate Centred Average of these 4 months moving Average.
  3. Then % Average ie Sales/Centred Average
  4. Then we table the % Average for each month (1-6) for the 2 years (2015-2016) and calculate mean for each month. (As Shown in Table 2)
  5. 600%/The total of these 6 months mean will be the x Adjustment Factor
  6. Seasonality Index for each month will be X Adj Factor * Month’s Mean ie Seasonality Index for Jan = 103.15%*1.01 = 1.04
  7. Then we calculate ∑X, ∑Y, ∑XY, ∑X2 and using these values we calculate the regression equation.
  8. Regression Equation will be: Y = 127.82 + 4.67*X
  9. Using this equation and Seasonality Index , we calculate value for Predicted Sales for next 6 months. Eg: Jan17 Sales = (127.82 + 4.67*13)*1.04

Table 1

Month

Month

X

Y

4 Q MA

Centred Average

% Average

XY

X2

SEASONAL INDEX

Jan-15

1

1

150

150

1

1.04

Feb-15

2

2

140

135.00

280

4

0.99

Mar-15

3

3

130

125.00

130.00

100.00%

390

9

0.99

Apr-15

4

4

120

137.50

131.25

91.43%

480

16

0.99

May-15

5

5

110

149.00

143.25

76.79%

550

25

0.77

Jun-15

6

6

190

162.75

155.88

121.89%

1140

36

1.23

Jan-16

1

7

176

178.50

170.63

103.15%

1232

49

1.04

Feb-16

2

8

175

177.25

177.88

98.38%

1400

64

0.99

Mar-16

3

9

173

179.00

178.13

97.12%

1557

81

0.99

Apr-16

4

10

185

174.75

176.88

104.59%

1850

100

0.99

May-16

5

11

183

2013

121

0.77

Jun-16

6

12

158

1896

144

1.23

Table 2

Month

2015

2016

Mean

X ADJ Factor

SEASONAL INDEX

Cumulative Indx

2017 Sales (Predicted)

Regression Equation: Y = a + bX

1

103.15%

103.15%

1.01

1.04

13

194.12

n

12

2

98.38%

98.38%

1.01

0.99

14

189.66

b

4.57

3

100.00%

97.12%

98.56%

1.01

0.99

15

194.53

a

127.82

4

91.43%

104.59%

98.01%

1.01

0.99

16

197.95

5

76.79%

76.79%

1.01

0.77

17

158.61

∑ XY

     12,938

6

121.89%

121.89%

1.01

1.23

18

257.37

∑ X2

           650

Total

596.79%

∑ X

              78

∑ Y

        1,890

Final Result

Month

SEASONAL INDEX

2017 Sales (Predicted)

Jan

1.04

194.12

Feb

0.99

189.66

Mar

0.99

194.53

Apr

0.99

197.95

May

0.77

158.61

Jun

1.23

257.37

Regression Equation:

Y = 127.82 + 4.67*X